# Unveiling the Role of Learning Rate Schedules via Functional Scaling Laws

### v1

@@ C.3 The g Functions
It holds for any $t>0$ that
$g_{a}(t)\leqslant\frac{1}{a},\quad g_{a}(t)\leqslant\frac{\Gamma(a)}{(2t)^{a}}$
@@
The proof only involves calculating the integral. Notice that
$g_{a}(t)\leqslant\int_{0}^{1}u^{a-1}e^{-2ut}\mathop{}\!\mathrm{d}u\leqslant\int_{0}^{1}u^{a-1}\mathop{}\!\mathrm{d}u=\frac{1}{a},$
and
$g_{a}(t)\leqslant\frac{1}{(2t)^{a}}\int_{0}^{\infty}u^{a-1}e^{-u}\mathop{}\!\mathrm{d}u=\frac{\Gamma(a)}{(2t)^{a}},$
@@
For $a>0$ and $t=\Omega(1)$, if $M^{\beta}>t$, we have $g_{a}(t)\eqsim\frac{1}{t^{a}}$. If $a>1$, we have $\int_{0}^{t}g_{a}(\tau)\mathop{}\!\mathrm{d}\tau\eqsim 1$.
@@
$g_{a}(t)$ $=\int_{M^{-\beta}}^{1}u^{a-1}e^{-2ut}\mathop{}\!\mathrm{d}u$
$=\frac{1}{(2t)^{a}}\int_{tM^{-\beta}}^{t}u^{a-1}e^{-u}\mathop{}\!\mathrm{d}u.$
The integral part is a partial gamma function, hence when $tM^{-\beta}\lesssim 1$, we have $g_{a}(t)\eqsim\frac{1}{t^{a}}$.
When $a>1$, we have
$\int_{0}^{t}g_{a}(\tau)\mathop{}\!\mathrm{d}\tau$ $=\int_{M^{-\beta}}^{1}\int_{0}^{t}u^{a-1}e^{-2u\tau}\mathop{}\!\mathrm{d}\tau\mathop{}\!\mathrm{d}u$
$=\int_{M^{-\beta}}^{1}u^{a-1}\frac{1-e^{-2ut}}{2u}\mathop{}\!\mathrm{d}u$
$=\frac{1-M^{-(a-1)\beta}}{2(a-1)}-\frac{1}{2}g_{a-1}(t)\lesssim 1.$
On the other hand, when $t=\Omega(1)$, $g_{a-1}(t)<\frac{1}{(a-1)}$, combining with $M\gg 1$ we have $\int_{0}^{t}g_{a}(\tau)\mathop{}\!\mathrm{d}\tau\eqsim 1$. ∎
@@
For $a,b\geqslant 0,t>0$, we have
$\int_{1}^{t-1}\frac{1}{(t-\tau)^{a}}\frac{1}{\tau^{b}}\mathop{}\!\mathrm{d}\tau\leqslant\frac{8}{t^{a\wedge b}},$
where $a\wedge b:=\min\{a,b\}$.
@@
Without loss of generality, we can assume $a\geqslant b$. By Hölder’s Inequality, we have
$\int_{1}^{t-1}\frac{1}{(t-\tau)^{a}}\frac{1}{\tau^{b}}\mathop{}\!\mathrm{d}\tau$ $=\int_{1}^{t-1}\frac{1}{(\tau(t-\tau))^{b}}\frac{1}{(t-\tau)^{a-b}}\mathop{}\!\mathrm{d}\tau$
$=\int_{1}^{t-1}\frac{1}{(\tau(t-\tau))^{b}}\frac{1}{(t-\tau)^{a-b}}\mathop{}\!\mathrm{d}\tau$
@@
$\leqslant\frac{2^{b-1}}{t^{b}}\int_{1}^{t-1}\left(\frac{1}{(t-\tau)^{a}}+\frac{1}{\tau^{b}(t-\tau)^{a-b}}\right)\mathop{}\!\mathrm{d}\tau$
$\leqslant\frac{2^{b}}{t^{b}}\int_{1}^{t-1}\frac{1}{\tau^{a}}\mathop{}\!\mathrm{d}\tau$
$\leqslant\frac{2^{b}a}{t^{b}}\leqslant\frac{8}{t^{b}}.$
@@
Assume $t,M$ is sufficiently large and $M^{\beta}>t$. For $a,b\geqslant 0,t>0$, we have
$\int_{0}^{t}g_{a}(t-\tau)g_{b}(\tau)\mathop{}\!\mathrm{d}\tau\eqsim g_{a\wedge b}(t),$
where $a\wedge b:=\min\{a,b\}$.
@@
Without loss of generality, we can assume $a\geqslant b$. Then, using Fubini’s theorem gives
$\int_{0}^{t}g_{a}(t-\tau)g_{b}(\tau)\mathop{}\!\mathrm{d}\tau$ $=\int_{0}^{t}\int_{M^{-\beta}}^{1}\int_{M^{-\beta}}^{1}u^{a-1}v^{b-1}e^{-2u(t-\tau)}e^{-2v\tau}\mathop{}\!\mathrm{d}u\mathop{}\!\mathrm{d}v\mathop{}\!\mathrm{d}\tau$
$=\int_{M^{-\beta}}^{1}\int_{M^{-\beta}}^{1}u^{a-1}v^{b-1}\frac{e^{-2vt}-e^{-2ut}}{2(u-v)}\mathop{}\!\mathrm{d}u\mathop{}\!\mathrm{d}v$
$=\int_{M^{-\beta}}^{1}v^{b-1}e^{-2vt}\left(\int_{M^{-\beta}}^{1}u^{a-1}\frac{1-e^{-2(u-v)t}}{2(u-v)}\mathop{}\!\mathrm{d}u\right)\mathop{}\!\mathrm{d}v$
Lower bound: Let $\epsilon_{1}=\frac{\log 2}{2t}$ and we have
$\int_{M^{-\beta}}^{1}u^{a-1}\frac{1-e^{-2(u-v)t}}{2(u-v)}\mathop{}\!\mathrm{d}u\geqslant\int_{M^{-\beta}}^{v-\epsilon}u^{a-1}\frac{1-e^{-2(u-v)t}}{2(u-v)}\mathop{}\!\mathrm{d}u+\int_{v+\epsilon}^{1}u^{a-1}\frac{1-e^{-2(u-v)t}}{2(u-v)}\mathop{}\!\mathrm{d}u.$
Notice that when $|u-v|\geqslant\frac{\log 2}{2t}$
$\frac{1-e^{-2(u-v)t}}{2(u-v)}\geqslant\frac{1}{4},$
and we have
$\int_{M^{-\beta}}^{1}u^{a-1}\frac{1-e^{-2(u-v)t}}{2(u-v)}\mathop{}\!\mathrm{d}u\geqslant\frac{1}{4}\left(\int_{M^{-\beta}}^{v-\epsilon}u^{a-1}\mathop{}\!\mathrm{d}u+\int_{v+\epsilon}^{1}u^{a-1}\mathop{}\!\mathrm{d}u\right)=\frac{1}{4}\left(1-M^{-a\beta}-\int_{v-\epsilon}^{v+\epsilon}u^{a-1}\mathop{}\!\mathrm{d}u\right).$
Notice that we have
$u^{a-1}\leqslant 1,$
and thus
$\int_{M^{-\beta}}^{1}u^{a-1}\frac{1-e^{-2(u-v)t}}{2(u-v)}\mathop{}\!\mathrm{d}u\geqslant\frac{1}{4}\left(1-M^{-a\beta}-\frac{\log 2}{t}\right).$
As $M,t=\Omega(1)$, we have
$M^{-a\beta}<\frac{1}{4},$
and
$\frac{\log 2}{t}<\frac{1}{4}.$
Combining them together, we have
$\int_{M^{-\beta}}^{1}u^{a-1}\frac{1-e^{-2(u-v)t}}{2(u-v)}\mathop{}\!\mathrm{d}u\geqslant\frac{1}{8},$
which implies that
$\int_{0}^{t}g_{a}(t-\tau)g_{b}(\tau)\mathop{}\!\mathrm{d}\tau\geqslant\frac{1}{8}g_{M}(a\wedge b,t).$
Upper bound: By lemma C.6
$\int_{0}^{t}g_{a}(t-\tau)g_{b}(\tau)\mathop{}\!\mathrm{d}\tau$ $\leqslant\frac{\Gamma(a)\Gamma(b)}{2^{a+b}}\int_{1}^{t-1}\frac{1}{(t-\tau)^{a}}\frac{1}{\tau^{b}}\mathop{}\!\mathrm{d}\tau$
$+\int_{0}^{1}g_{a}(t-\tau)g_{b}(\tau)\mathop{}\!\mathrm{d}\tau+\int_{t-1}^{t}g_{a}(t-\tau)g_{b}(\tau)\mathop{}\!\mathrm{d}\tau$
We observe that
$\int_{t-1}^{t}g_{a}(t-\tau)g_{b}(\tau)\mathop{}\!\mathrm{d}\tau\eqsim\frac{1}{a}\int_{t-1}^{t}g_{b}(\tau)\mathop{}\!\mathrm{d}\tau\eqsim\frac{1}{a}\frac{1}{t^{b}}.$
and for the same reason
$\int_{0}^{1}g_{a}(t-\tau)g_{b}(\tau)\mathop{}\!\mathrm{d}\tau\eqsim\frac{1}{b}\int_{0}^{1}g_{a}(\tau)\mathop{}\!\mathrm{d}\tau\eqsim\frac{1}{b}\frac{1}{t^{a}}\leqslant\frac{1}{b}\frac{1}{t^{b}}.$
On the other hand, by the Lemma C.8 we have
$\int_{1}^{t-1}\frac{1}{(t-\tau)^{a}}\frac{1}{\tau^{b}}\mathop{}\!\mathrm{d}\tau\leqslant\frac{8}{t^{a\wedge b}}.$
Combine them together we have
$\int_{0}^{t}g_{a}(t-\tau)g_{b}(\tau)\mathop{}\!\mathrm{d}\tau\leqslant\left(\frac{1}{a}+\frac{1}{b}+8\right)\frac{1}{t^{b}}\leqslant\frac{10}{t^{b}}.$
Observe that $M^{-\beta}t<1$ and $t>t_{0}$, we have
$g_{b}(t)=\frac{1}{t^{b}}\int_{M^{-\beta}t}^{t}u^{b-1}e^{-2u}\mathop{}\!\mathrm{d}u\geqslant\frac{1}{t^{b}}\int_{1}^{t_{0}}u^{b-1}e^{-2u}\mathop{}\!\mathrm{d}u=\frac{C_{1}}{t^{b}}$
So we can give the upper bound
$\int_{0}^{t}g_{a}(t-\tau)g_{b}(\tau)\mathop{}\!\mathrm{d}\tau\leqslant 10\frac{1}{t^{b}}\leqslant\frac{10}{C_{1}}g_{a\wedge b}(t).$
By combining the lower and upper bounds, we complete the proof of the lemma. ∎
Recall the definition of $G_{a}(\cdot)$. Then, the following lemma follows from Lemma C.9 immediately by taking $M\to\infty$.

### v3

@@ C.2 Power-law Decay of the $e_{M}$ and $\mathcal{K}_{M}$ Functions
For $e_{M}(t)$, $\mathcal{K}_{M}(t)$ we have the following estimation:
$e_{M}(t)\lesssim t^{-s},\quad\mathcal{K}_{M}(t)\lesssim t^{-(2-\frac{1}{\beta})}.$
Moreover,
$e_{M}(t)\eqsim t^{-s},\quad\mathcal{K}_{M}(t)\eqsim t^{-(2-\frac{1}{\beta})},\quad 1\lesssim t\lesssim M^{\beta}.$
@@
Recall that
$e_{M}(t)=\sum_{j=1}^{M}e^{-2\lambda_{j}t}\lambda_{j}|\theta_{j}^{*}|^{2}\eqsim\sum_{j=1}^{M}e^{-2\lambda_{j}t}j^{-1-s\beta}.$
Since $\lambda_{j}\eqsim j^{-\beta}$, there exists a constant $c$ such that
$\lambda_{j}\leqslant cj^{-\beta}\implies 2\lambda_{j}t\leqslant 1\text{ when }j\geqslant(2ct)^{\frac{1}{\beta}}.$
We have
$e_{M}(t)$ $\eqsim\sum_{j=1}^{[(2ct)^{\frac{1}{\beta}}]}e^{-2\lambda_{j}t}j^{-1-s\beta}+\sum_{j=[(2ct)^{\frac{1}{\beta}}]+1}^{M}e^{-2\lambda_{j}t}j^{-1-s\beta}$
$\lesssim(2ct)^{\frac{1}{\beta}}\max_{j}e^{-2\lambda_{j}t}j^{-1-s\beta}+\sum_{j=[(2ct)^{\frac{1}{\beta}}]+1}^{M}j^{-1-s\beta}.$
Now we bound the two terms separately:
$\max_{j}e^{-2\lambda_{j}t}j^{-1-s\beta}\eqsim\max_{j}e^{-2\lambda_{j}t}(\lambda_{j}t)^{s+\frac{1}{\beta}}\cdot t^{-s-\frac{1}{\beta}}\leqslant t^{-s-\frac{1}{\beta}}\sup_{x>0}e^{-2x}x^{s+\frac{1}{\beta}}\lesssim t^{-s-\frac{1}{\beta}}.$
where the last inequality is because $\sup_{x>0}e^{-2x}x^{s+\frac{1}{\beta}}$ is a constant only depending on $s$ and $\beta$.
We have
$\sum_{j=[(2ct)^{\frac{1}{\beta}}]+1}^{M}j^{-1-s\beta}\lesssim\int_{[(2ct)^{\frac{1}{\beta}}]}^{\infty}x^{-1-s\beta}\mathop{}\!\mathrm{d}x\eqsim[(2ct)^{\frac{1}{\beta}}]^{-s\beta}\eqsim t^{-s}.$
Therefore we have for all $t>0$,
$e_{M}(t)\lesssim(2ct)^{\frac{1}{\beta}}t^{-s-\frac{1}{\beta}}+t^{-s}\eqsim t^{-s}.$
On the other hand, when $M>2([(2ct)^{\frac{1}{\beta}}+1])$, i.e., $t\lesssim M^{\beta}$,
$e_{M}(t)$ $\eqsim\sum_{j=1}^{[(2ct)^{\frac{1}{\beta}}]}e^{-2\lambda_{j}t}j^{-1-s\beta}+\sum_{j=[(2ct)^{\frac{1}{\beta}}]+1}^{M}e^{-2\lambda_{j}t}j^{-1-s\beta}$
$\gtrsim e^{-1}\sum_{j=[(2ct)^{\frac{1}{\beta}}]+1}^{M}j^{-1-s\beta}$
$\gtrsim\int_{[(2ct)^{\frac{1}{\beta}}]+1}^{2[(2ct)^{\frac{1}{\beta}}]+1}x^{-1-s\beta}\mathop{}\!\mathrm{d}x$
$=\frac{1}{s\beta}(([(2ct)^{\frac{1}{\beta}}+1])^{-s\beta}-(2[(2ct)^{\frac{1}{\beta}}]+1)^{-s\beta})$
$\eqsim t^{-s},$
where the last line requires $t\gtrsim 1$. Therefore we have shown that $e_{M}(t)\eqsim t^{-s}$ when $1\lesssim t\lesssim M^{\beta}$.
The proof for $\mathcal{K}_{M}(t)$ follows the same way. ∎
@@
We can bound the gap between $e_{M}(t)$ and $e_{\infty}(t)$ as follows:
$e_{\infty}(t)-e_{M}(t)\lesssim M^{-s\beta}.$
As a consequence, we have
$e_{M}(t)+M^{-s\beta}\eqsim e_{\infty}(t)+M^{-s\beta}.$
@@
$e_{\infty}(t)-e_{M}(t)$ $=\sum_{j=M+1}^{\infty}e^{-2\lambda_{j}t}\lambda_{j}|\theta_{j}^{*}|^{2}$
$\lesssim\sum_{j=M+1}^{\infty}j^{-1-s\beta}$
$\leqslant\int_{M}^{\infty}x^{-1-s\beta}\mathop{}\!\mathrm{d}x\eqsim M^{-s\beta}.$
C.3 The Case of Top-$M$ Features
First, we prove the general label-noise case of FSL (Theorem 4.3). Applying this result, we then derive the constant label-noise case (Theorem 4.4), the noiseless case (Theorem 4.5) and the hard regime case (Theorem 4.2)
C.3.1 Proof of Theorem 4.3
In the top-$M$ feature case, the matrix $\mathbf{W}$ satisfies $\mathbf{w}_{j}=\mathrm{\mathbf{e}}_{j}$ for each $j\in[M]$, therefore we can simplify the equation for $\mathbb{E}[\mathcal{E}_{t}]$ as follows.
Theorem C.6 (Volterra equation of the top-$M$ case).
In the top-$M$ case, we have
$\mathbb{E}[\mathcal{E}_{t}]\eqsim M^{-s\beta}+e_{M}(t)+\int_{0}^{t}\mathcal{K}_{M}(t-\tau)\gamma_{\tau}(\mathbb{E}[\mathcal{E}_{\tau}]+\sigma^{2})\mathop{}\!\mathrm{d}\tau,$ (35)
where the function $e_{M}$ and $\mathcal{K}_{M}$ are defined as
$e_{M}(t):=\sum_{j=1}^{M}\lambda_{j}(\theta_{j}^{*})^{2}e^{-2\lambda_{j}t},\quad\mathcal{K}_{M}(t):=\sum_{j=1}^{M}\lambda_{j}^{2}e^{-2\lambda_{j}t}.$ (36)
Proof.
By (34) in Proposition C.3, note that $\mathbf{W}^{\top}\mathbf{W}\mathbf{H}=\mathbf{H}_{0:M}\in\mathbb{R}^{N\times N}$ is the top-$M$ part of the matrix $\mathbf{H}$, i.e. $\mathbf{H}_{0:M}=\operatorname{diag}\{\lambda_{1},\dots,\lambda_{M},0,\dots,0\}$, we get
$\mathbb{E}[\mathcal{E}_{t}]\eqsim\mathbf{u}_{0}^{\top}\mathbf{H}e^{-2\mathbf{H}_{0:M}t}\mathbf{u}_{0}+\int_{0}^{t}\mathrm{tr}(\mathbf{H}_{0:M}^{2}e^{-2\mathbf{H}_{0:M}})\gamma_{\tau}(\mathbb{E}[\mathcal{E}_{t}]+\sigma^{2})\mathop{}\!\mathrm{d}\tau,$
which can be further written in terms of the eigenvalues $\{\lambda_{j}\}$ as
$\mathbb{E}[\mathcal{E}_{t}]\eqsim\sum_{j=1}^{M}\lambda_{j}(u_{0}^{(j)})^{2}e^{-2\lambda_{j}t}+\sum_{j=M+1}^{\infty}\lambda_{j}(u_{0}^{(j)})^{2}+\int_{0}^{t}\sum_{j=1}^{M}\lambda_{j}^{2}e^{-2\lambda_{j}t}\gamma_{\tau}(\mathbb{E}[\mathcal{E}_{t}]+\sigma^{2})\mathop{}\!\mathrm{d}\tau.$
Note that $u_{0}^{(j)}$, the $j$-th component of $\mathbf{u}_{0}$, is equal to $\theta_{j}^{*}$ because of the zero initialization of $\bm{\nu}_{0}$.
Therefore by the definition of $e_{M}$ and $\mathcal{K}_{M}$ we arrive at the Volterra-type integral equation of $\mathbb{E}[\mathcal{E}_{t}]$. ∎
@@ C.3 The Case of Top-$M$ Features
For the forgetting kernel $\mathcal{K}_{M}$ and $t\leqslant M^{\beta}$, there exists a constant $C$ independent of $t$, such that
$\mathcal{K}_{M}*\mathcal{K}_{M}(t)\leqslant C\mathcal{K}_{M}(t),\quad\forall t\leqslant M^{\beta}.$
where * denotes convolution:
$\mathcal{K}_{M}*\mathcal{K}_{M}(t):=\int_{0}^{t}\mathcal{K}_{M}(\tau)\mathcal{K}_{M}(t-\tau)\mathop{}\!\mathrm{d}\tau.$
@@
Observe that $\mathcal{K}_{M}(t)$ is a monotonically decreasing function. By the symmetry of the convolution, we can write
$\mathcal{K}_{M}*\mathcal{K}_{M}(t)$ $=\int_{0}^{t}\mathcal{K}_{M}(\tau)\mathcal{K}_{M}(t-\tau)\,\mathop{}\!\mathrm{d}\tau$
$=2\int_{0}^{t/2}\mathcal{K}_{M}(\tau)\mathcal{K}_{M}(t-\tau)\,\mathop{}\!\mathrm{d}\tau.$
Since $\mathcal{K}_{M}$ is decreasing, for $0\leqslant\tau\leqslant t/2$ we have $\mathcal{K}_{M}(t-\tau)\leqslant\mathcal{K}_{M}(t/2)$. This observation yields the upper bound
@@
$\mathcal{K}_{M}*\mathcal{K}_{M}(t)\leqslant 2\,\mathcal{K}_{M}\!\left(\tfrac{t}{2}\right)\int_{0}^{\infty}\mathcal{K}_{M}(\tau)\,\mathop{}\!\mathrm{d}\tau\leqslant C\,\mathcal{K}_{M}\!\left(\tfrac{t}{2}\right),$
where the constant $C$ is finite and given by
$C$ $=2\int_{0}^{\infty}\mathcal{K}_{M}(\tau)\,\mathop{}\!\mathrm{d}\tau\leqslant 2\int_{0}^{\infty}\sum_{j=1}^{\infty}\lambda_{j}^{2}e^{-2\lambda_{j}\tau}\,\mathop{}\!\mathrm{d}\tau=\sum_{j=1}^{\infty}\lambda_{j}=\mathrm{tr}(\mathbf{H})<\infty.$
It remains to show that when $t\leqslant M^{\beta}$,
$\mathcal{K}_{M}\left(\frac{t}{2}\right)\leqslant C^{\prime}\mathcal{K}_{M}(t)$
for some constant $C^{\prime}>0$. Recall that by Lemma C.4 we have
$\mathcal{K}_{M}(t)\eqsim t^{-(2-\frac{1}{\beta})},\quad\mathcal{K}_{M}\left(\frac{t}{2}\right)\eqsim\left(\frac{t}{2}\right)^{-(2-\frac{1}{\beta})}.$
Therefore when $1\lesssim t\lesssim M^{\beta}$ the conclusion follows. When $t\lesssim 1$, the function $\mathcal{K}_{M}(t)$ is decreasing and $\mathcal{K}_{M}(t)\leqslant\sum_{j=1}^{M}\lambda_{j}^{2}<\infty$, we can always find a constant $C^{\prime}$ satisfying the condition.
Combining the above estimates, we obtain
$(\mathcal{K}_{M}*\mathcal{K}_{M})(t)\leqslant C\,\mathcal{K}_{M}\left(\tfrac{t}{2}\right)\leqslant CC_{1}\,\mathcal{K}_{M}(t)=:C^{\prime}\mathcal{K}_{M}(t).$
The proof is complete. ∎
We now prove Theorem 4.3.
Proof.
The lower bound is trivial by $\mathbb{E}[\mathcal{E}_{t}]\geqslant e_{M}(t)$ and the Volterra equation (35):
$\mathbb{E}[\mathcal{E}_{t}]$ $\eqsim M^{-s\beta}+e_{M}(t)+\int_{0}^{t}\mathcal{K}_{M}(t-\tau)\gamma_{\tau}(\mathbb{E}[\mathcal{E}_{\tau}]+\sigma^{2})\mathop{}\!\mathrm{d}\tau$
$\gtrsim M^{-s\beta}+e_{M}(t)+\int_{0}^{t}\mathcal{K}_{M}(t-\tau)\gamma_{\tau}(e_{M}(\tau)+\sigma^{2})\mathop{}\!\mathrm{d}\tau.$
For the upper bounds, we first prove a slightly stronger bound with $\mathcal{K}_{\infty}$, that is,
Proposition C.8.
For all $t>0$, $s>0$ and $\beta>1$, we have
$\mathbb{E}[\mathcal{E}_{t}]\lesssim M^{-s\beta}+e_{M}(t)+\int_{0}^{t}\mathcal{K}_{\infty}(t-\tau)(e_{M}(\tau)+\sigma^{2})\gamma_{\tau}\mathop{}\!\mathrm{d}\tau.$ (37)
Proof of C.8.
By Equation (35), we have
$\mathbb{E}[\mathcal{E}_{t}]$ $\eqsim M^{-s\beta}+e_{M}(t)+\int_{0}^{t}\mathcal{K}_{M}(t-\tau)\gamma_{\tau}(\mathbb{E}[\mathcal{E}_{\tau}]+\sigma^{2})\mathop{}\!\mathrm{d}\tau$
$\leqslant M^{-s\beta}+e_{M}(t)+\int_{0}^{t}\mathcal{K}_{\infty}(t-\tau)\gamma_{\tau}(\mathbb{E}[\mathcal{E}_{\tau}]+\sigma^{2})\mathop{}\!\mathrm{d}\tau.$
Define $f(t):=\int_{0}^{t}\mathcal{K}_{\infty}(t-\tau)\gamma_{\tau}(\mathbb{E}[\mathcal{E}_{\tau}]+\sigma^{2})\mathop{}\!\mathrm{d}\tau$ and $\gamma_{\max}=\sup_{t}\gamma(t)$, substituting the above inequality into the definition of $f(t)$, we get
$f(t)$ $\lesssim\int_{0}^{t}\mathcal{K}_{\infty}(t-\tau)(M^{-s\beta}+e_{M}(\tau)+\sigma^{2})\gamma_{\tau}\mathop{}\!\mathrm{d}\tau$
$\quad+\int_{0}^{t}\int_{0}^{\tau}\mathcal{K}_{\infty}(t-\tau)\mathcal{K}_{\infty}(\tau-r)\gamma_{\tau}\gamma_{r}(\mathbb{E}[\mathcal{E}_{r}]+\sigma^{2})\mathop{}\!\mathrm{d}r\mathop{}\!\mathrm{d}\tau$
$\lesssim\gamma_{\max}M^{-s\beta}+\int_{0}^{t}\mathcal{K}_{\infty}(t-\tau)(e_{M}(\tau)+\sigma^{2})\gamma_{\tau}\mathop{}\!\mathrm{d}\tau$
$\quad+\gamma_{\max}\int_{0}^{t}\int_{r}^{t}\mathcal{K}_{\infty}(t-\tau)\mathcal{K}_{\infty}(\tau-r)\mathop{}\!\mathrm{d}\tau\gamma_{r}(\mathbb{E}[\mathcal{E}_{r}]+\sigma^{2})\mathop{}\!\mathrm{d}r$
$=\gamma_{\max}M^{-s\beta}+\int_{0}^{t}\mathcal{K}_{\infty}(t-\tau)(e_{M}(\tau)+\sigma^{2})\gamma_{\tau}\mathop{}\!\mathrm{d}\tau$
$\quad+\gamma_{\max}\int_{0}^{t}(\mathcal{K}_{\infty}*\mathcal{K}_{\infty})(t-r)\gamma_{r}(\mathbb{E}[\mathcal{E}_{r}]+\sigma^{2})\mathop{}\!\mathrm{d}r$
[…]
Therefore when $\gamma_{\max}$ is sufficiently small ($\gamma_{\max}\leqslant\frac{c}{\mathrm{tr}(\mathbf{H})}$ for some absolute constant $c$), the constant factor of $f(t)$ on the right-hand side will be less than $\frac{1}{2}$, hence we may subtract $\gamma_{\max}f(t)$ from both sides and get
$\int_{0}^{t}\mathcal{K}_{\infty}(t-\tau)\gamma_{\tau}\mathbb{E}[\mathcal{E}_{\tau}]\mathop{}\!\mathrm{d}\tau\lesssim\gamma_{\max}M^{-s\beta}+\int_{0}^{t}\mathcal{K}_{\infty}(t-\tau)(e_{M}(\tau)+\sigma^{2})\gamma_{\tau}\mathop{}\!\mathrm{d}\tau.$
Therefore substituting this back to the Volterra equation yields
$\mathbb{E}[\mathcal{E}_{t}]\lesssim M^{-s\beta}+e_{M}(t)+\int_{0}^{t}\mathcal{K}_{\infty}(t-\tau)(e_{M}(\tau)+\sigma^{2})\gamma_{\tau}\mathop{}\!\mathrm{d}\tau.$
The preceding proposition relies exclusively on two properties of $\mathcal{K}_{\infty}$: the identity established in Lemma C.7 and the finiteness of its integral, $\int_{0}^{\infty}\mathcal{K}_{\infty}(t)\mathop{}\!\mathrm{d}t<\infty$. Since the kernel $\mathcal{K}_{M}$ also possesses these properties for $t\leqslant M^{\beta}$, the derivation of the upper bound
$\mathbb{E}[\mathcal{E}_{t}]\lesssim M^{-s\beta}+e_{M}(t)+\int_{0}^{t}\mathcal{K}_{M}(t-\tau)(e_{M}(\tau)+\sigma^{2})\gamma_{\tau}\mathop{}\!\mathrm{d}\tau,\quad t\leqslant M^{\beta}$
follows analogously by substituting $\mathcal{K}_{\infty}$ with $\mathcal{K}_{M}$. ∎
C.3.2 Proof of Theorem 4.4
Proof.
It is clear that, by our assumption, $\sup_{t}\gamma_{t}\lesssim 1$, and that
$e_{M}(t)\eqsim\sum_{j=1}^{M}j^{-1-s\beta}e^{-2\lambda_{j}t}\leqslant\sum_{j=1}^{M}j^{-1-s\beta}\eqsim 1.$
By Proposition C.8, we have
$\mathbb{E}[\mathcal{E}_{t}]$ $\lesssim M^{-s\beta}+e_{M}(t)+\int_{0}^{t}\mathcal{K}_{\infty}(t-\tau)(e_{M}(t)+\sigma^{2})\gamma_{\tau}\mathop{}\!\mathrm{d}\tau$ (38)
$\lesssim 1+\int_{0}^{t}\mathcal{K}_{\infty}(t-\tau)(\sigma^{2}+1)\mathop{}\!\mathrm{d}\tau\lesssim\sigma^{2},$ (39)
where the last inequality is by $\sigma^{2}\gtrsim 1$ and that
$\int_{0}^{t}\mathcal{K}_{\infty}(\tau)\mathop{}\!\mathrm{d}\tau=\int_{0}^{t}\sum_{j=1}^{\infty}\lambda_{j}^{2}e^{-2\lambda_{j}\tau}\mathop{}\!\mathrm{d}\tau=\frac{1}{2}\sum_{j=1}^{\infty}\lambda_{j}\eqsim 1.$
Therefore by the Volterra equation (35), note that $\mathbb{E}[\mathcal{E}_{\tau}]+\sigma^{2}\eqsim\sigma^{2}\eqsim e_{M}(\tau)+\sigma^{2}$, we have
$\mathbb{E}[\mathcal{E}_{t}]\eqsim M^{-s\beta}+e_{M}(t)+\int_{0}^{t}\mathcal{K}_{M}(t-\tau)(e_{M}(\tau)+\sigma^{2})\gamma_{\tau}\mathop{}\!\mathrm{d}\tau.$
C.3.3 Proof of Theorem 4.5
Proof.
When $\sigma^{2}=0$, by FSL in general case (Proposition C.8), we have
$\mathbb{E}[\mathcal{E}_{t}]\lesssim M^{-s\beta}+e_{M}(t)+\int_{0}^{t}\mathcal{K}_{\infty}(t-\tau)e_{M}(\tau)\gamma_{\tau}\mathop{}\!\mathrm{d}\tau.$
In order to get the upper bound with $\mathcal{K}_{M}$, we first prove a lemma.
@@
We will bound the gap introduced by $\mathcal{K}_{\infty}$ and $\mathcal{K}_{M}$:
$\int_{0}^{t}(\mathcal{K}_{\infty}(t-\tau)-\mathcal{K}_{M}(t-\tau))e_{M}(\tau)\gamma_{\tau}\mathop{}\!\mathrm{d}\tau\lesssim M^{\max\{-s\beta,-2\beta+1\}}.$ (40)
Proof of Lemma.
First note that
$\mathcal{K}_{\infty}(t)-\mathcal{K}_{M}(t)=\sum_{j=M+1}^{\infty}\lambda_{j}^{2}e^{-2\lambda_{j}t}\lesssim\sum_{j=M+1}^{\infty}j^{-2\beta}\eqsim M^{-2\beta+1}.$
Therefore we can bound the integral as
$\int_{0}^{t}(\mathcal{K}_{\infty}(t-\tau)-\mathcal{K}_{M}(t-\tau))e_{M}(\tau)\gamma_{\tau}\mathop{}\!\mathrm{d}\tau$
$\lesssim M^{-2\beta+1}\int_{0}^{t}e_{M}(\tau)\gamma_{\tau}\mathop{}\!\mathrm{d}\tau$
$\leqslant\gamma_{\max}M^{-2\beta+1}\int_{0}^{\infty}e_{M}(\tau)\mathop{}\!\mathrm{d}\tau$
$=\gamma_{\max}M^{-2\beta+1}\int_{0}^{\infty}\sum_{j=1}^{M}\lambda_{j}(\theta_{j}^{*})^{2}e^{-2\lambda_{j}\tau}\mathop{}\!\mathrm{d}\tau$
$\eqsim\gamma_{\max}M^{-2\beta+1}\sum_{j=1}^{M}j^{-s\beta-1+\beta}$
$\lesssim\gamma_{\max}M^{\max\{-s\beta-\beta+1,-2\beta+1\}}\lesssim\gamma_{\max}M^{\max\{-s\beta,-2\beta+1\}}$
By $s\leqslant 2-\frac{1}{\beta}$ and $\gamma_{\max}$ is sufficiently small, we can combine the upper bound with (40), and directly conclude that
$\mathbb{E}[\mathcal{E}_{t}]\lesssim M^{-s\beta}+e_{M}(t)+\int_{0}^{t}\mathcal{K}_{M}(t-\tau)e_{M}(\tau)\gamma_{\tau}\mathop{}\!\mathrm{d}\tau.$
Now with the lower bound in FSL Theorem 4.3, the result follows. ∎
C.3.4 Proof of Theorem 4.2
@@
In the hard regime $s\leqslant 1-\frac{1}{\beta}$, by Propostion C.8, we have
$\mathbb{E}[\mathcal{E}_{t}]\lesssim M^{-s\beta}+e_{M}(t)+\int_{0}^{t}\mathcal{K}_{\infty}(t-\tau)(e_{M}(\tau)+\sigma^{2})\gamma_{\tau}\mathop{}\!\mathrm{d}\tau.$